# Forecasting realized volatility in the stock market: a path-dependent perspective

### v1

@@ 4.5.1 Out-of-sample $R^{2}$ test
Here, $\mathrm{RV}_{t}$ represents the actual realized volatility, $\mathrm{RV}^{j}_{t}$ denotes the out-of-sample predicted value from the forecasting model, and $\mathrm{RV}^{0}_{t}$ refers to the out-of-sample predicted value from the benchmark model. If $R^{2}_{\mathrm{oos}}(\%)$ 0, it indicates that the model’s mean squared predictive error (MSPE-adjusted) is lower than that of the benchmark model, demonstrating positive predictive capability. If $R^{2}_{\mathrm{oos}}(\%)$ = 0, the forecasting performance of the model is identical to that of the benchmark model; if $R^{2}_{\mathrm{oos}}(\%)<0$ , the model performs worse than the benchmark model. The method for calculating the adjusted MSPE is detailed in Clark and West (2007).
@@ 4.5.2 Different rolling window lengths
GARCH $-0.172$ $3.881\times 10^{-9}$ $0.265$
EGARCH $-0.633$ $-2.571\times 10^{-9}$ $0.852$
FGARCH $-0.625$ $-8.033\times 10^{-9}$ $0.868$
TGARCH $-0.164$ $4.756\times 10^{-9}$ $0.091$
@@
LASSO-HAR-PD-RS $0.263$ $2.632\times 10^{-8}$ $0.061$
LASSO-HAR-PD-REX $0.256$ $2.732\times 10^{-8}$ $0.065$
LASSO-HAR-PD-REQ $0.291$ $3.331\times 10^{-8}$ $0.102$
@@
As shown in Table 9, all models in the HAR-PD family exhibit an $R^{2}_{\mathrm{oos}}(\%)$ value significantly greater than zero, indicating their strong effectiveness in out-of-sample forecasting. Specifically, when comparing the HAR and HAR-PD model families, the HAR-PD-REX model demonstrates the highest $R^{2}_{\mathrm{oos}}(\%)$ value in out-of-sample forecasting.
Furthermore, in Table 9, the $R^{2}_{\mathrm{oos}}(\%)$ values of the LASSO-HAR-PD model family and HAR-PD model family are also higher than those of the HAR model family, highlighting the overall improvement in predictive accuracy of the HAR-PD model family over traditional HAR model family. These findings further confirm that volatility forecasting models constructed from a path-dependence perspective exhibit superior predictive accuracy compared to baseline models.
@@ 4.5.3 Different sample data
Table 10: Out-of-sample forecast loss value of CSI 300 for the future 1-step
@@
GARCH $1.117\times 10^{-8}$ $5.326\times 10^{-5}$ $1.922$ $0.853$ $-8.441$
EGARCH $2.444\times 10^{-8}$ $1.216\times 10^{-4}$ $11.766$ $2.556$ $-8.115$
FGARCH $1.701\times 10^{-8}$ $7.235\times 10^{-5}$ $5.397$ $1.361$ $-8.414$
TGARCH $1.478\times 10^{-8}$ $7.028\times 10^{-5}$ $4.267$ $1.277$ $-7.865$
HAR-RV $1.228\times 10^{-8}$ $5.466\times 10^{-5}$ $3.409$ $0.865$ $-8.364$
HAR-CJ $8.565\times 10^{-9}$ $3.831\times 10^{-5}$ $0.477$ 0.4960.496{0.496}0.496 $-8.536$
HAR-RS $7.919\times 10^{-9}$ $3.905\times 10^{-5}$ $1.003$ $0.571$ $-8.553$
HAR-REX $9.162\times 10^{-9}$ $4.421\times 10^{-5}$ $0.800$ $0.666$ $-8.513$
HAR-REQ $8.866\times 10^{-9}$ ${3.818\times 10^{-5}}$ $\underline{\mathbf{0.433}}$ $\underline{\mathbf{0.484}}$ $-8.525$
HAR-PD-RV $9.513\times 10^{-9}$ $4.577\times 10^{-5}$ $1.241$ $0.731$ $-8.501$
hHAR-PD-CJ $8.422\times 10^{-9}$ $4.226\times 10^{-5}$ $0.688$ $0.604$ $-8.529$
HAR-PD-RS $7.774\times 10^{-9}$ $3.886\times 10^{-5}$ $1.771$ $0.574$ $\underline{\mathbf{-8.555}}$
HAR-PD-REX $8.064\times 10^{-9}$ $4.062\times 10^{-5}$ $0.640$ $0.593$ $-8.535$
HAR-PD-REQ ${7.740\times 10^{-9}}$ $3.956\times 10^{-5}$ $0.602$ $0.569$ $-8.554$
LASSO-HAR-PD-CJ $8.427\times 10^{-9}$ $4.250\times 10^{-5}$ $0.703$ $0.612$ $-8.528$
LASSO-HAR-PD-RS $\underline{\mathbf{7.638\times 10^{-9}}}$ $\underline{\mathbf{3.757\times 10^{-5}}}$ $0.956$ $0.590$ $-8.597$
LASSO-HAR-PD-REX $7.741\times 10^{-9}$ $3.951\times 10^{-5}$ $0.599$ $0.567$ $-8.553$
LASSSO-HAR-PD-REQ $8.283\times 10^{-9}$ $4.140\times 10^{-5}$ $0.647$ $0.600$ $-8.526$
@@
In Table 10, the loss function values for the future 1-step predictions indicate that the HAR-REQ and HAR-PD-REQ models exhibit lower loss values across two different loss functions, while the HAR-PD-RS model achieves the lowest loss value under the QLIKE loss function. Notably, the proposed HAR-REQ model and its improved HAR-PD-REQ variant outperform other models in volatility forecasting under the MAE, HMSE, and HMAE loss functions.
@@ 5 Conclusion
In out-of-sample forecasting, the HAR-PD model family, constructed based on path dependence, consistently outperformed the original HAR model family, significantly improving prediction accuracy. This result remained significant across different datasets, indicating the strong robustness of the new model. Notably, under the rolling window and out-of-sample evaluation methods, the new model exhibited sustained superiority. It is worth mentioning that among all models examined in this study, the HAR-PD-RS model achieved the highest prediction accuracy in forecasting daily volatility, while the HAR-PD-REQ model demonstrated superior predictive capability in medium- and long-term (5-step and 22-step) volatility forecasting.
@@ Appendix B Out-of-sample forecast loss value and MCS test
Table 11: MCS test for the future 1-step volatility forecasting model of the SSE
@@
GARCH $0.000$ $0.000$ $0.019$ $0.045$ $0.034$ $0.029$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
FGARCH $0.000$ $0.000$ $0.964$ $0.029$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
TGARCH $0.000$ $0.000$ $0.013$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-RV $0.000$ $0.000$ $0.235$ $0.120$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-CJ $0.000$ $0.000$ $0.183$ $0.525$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-RS $1.000$ $0.000$ $1.000$ $0.465$ $1.000$ $0.000$ $1.000$ $0.324$ $1.000$ $0.000$ $0.944$ $0.319$
HAR-REX $0.075$ $0.525$ $1.000$ $0.574$ $0.000$ $0.000$ $1.000$ $0.423$ $0.000$ $0.000$ $0.958$ $0.412$
HAR-REQ $0.391$ $0.679$ $1.000$ $1.000$ $1.000$ $1.000$ $0.243$ $0.000$ $0.000$ $0.000$ $1.000$ $1.000$
HAR-PD-RV $0.000$ $0.000$ $1.000$ $0.485$ $0.000$ $0.000$ $0.102$ $0.099$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-PD-CJ $0.712$ $0.705$ $1.000$ $1.000$ $0.000$ $1.000$ $0.802$ $1.000$ $1.000$ $0.827$ $0.499$ $1.000$
HAR-PD-RS $0.984$ $0.005$ $1.000$ $1.000$ $0.470$ $0.248$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$
HAR-PD-REX $1.000$ $0.994$ $1.000$ $\mathbf{1.000}$ $0.661$ $0.436$ $\mathbf{1.000}$ $\mathbf{1.000}$ $0.000$ $0.000$ $\mathbf{1.000}$ $\mathbf{1.000}$
HAR-PD-REQ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $0.000$ $0.000$ $1.000$ $1.000$
LASSO-HAR-PD-CJ $0.657$ $0.679$ $1.000$ $1.000$ $0.274$ $1.000$ $0.971$ $0.440$ $1.000$ $1.000$ $0.998$ $0.977$
@@
Table 12: MCS test for the future 5-step volatility forecasting model of the SSE
@@
FGARCH $0.000$ $0.000$ $0.964$ $0.029$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
TGARCH $0.000$ $0.000$ $0.047$ $0.067$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-CJ $0.227$ $0.996$ $1.000$ $0.453$ $0.000$ $0.000$ $1.000$ $0.423$ $0.000$ $0.000$ $0.993$ $0.408$
HAR-RS $1.000$ $0.006$ $1.000$ $0.095$ $1.000$ $0.000$ $1.000$ $0.085$ $1.000$ $0.000$ $0.906$ $0.089$
HAR-REX $0.138$ $0.622$ $1.000$ $0.946$ $0.000$ $0.000$ $1.000$ $0.423$ $0.000$ $0.000$ $0.868$ $0.462$
HAR-REQ $0.656$ $1.000$ $0.246$ $0.453$ $0.237$ $0.640$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$
HAR-PD-RV $0.000$ $0.000$ $1.000$ $0.485$ $0.000$ $0.000$ $0.107$ $0.097$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-PD-CJ $0.867$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $0.802$ $1.000$ $0.339$ $0.573$ $1.000$ $1.000$
HAR-PD-RS $0.946$ $0.004$ $1.000$ $1.000$ $0.665$ $0.640$ $1.000$ $1.000$ $0.557$ $0.512$ $1.000$ $1.000$
HAR-PD-REX $1.000$ $1.000$ $1.000$ 1.000 $1.000$ 1.000 $1.000$ 1.000 $0.845$ $0.983$ 1.000 1.000
HAR-PD-REQ $1.000$ $1.000$ 1.000 $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $0.000$ $1.000$ $1.000$
LASSO-HAR-PD-CJ $0.769$ $1.000$ $1.000$ $1.000$ $0.345$ $0.995$ $0.971$ $0.440$ $0.263$ $0.512$ $0.995$ $1.000$
@@
Table 13: MCS test for the future 22-step volatility forecasting model of the SSE
@@
Table 14: Out-of-sample forecast loss value of SSE for the future 5-step
@@
GARCH $1.636\times 10^{-8}$ $7.935\times 10^{-5}$ $10.693$ $1.872$ $-8.360$
EGARCH $1.174\times 10^{-8}$ $6.927\times 10^{-5}$ $5.059$ $1.511$ $-8.499$
FGARCH $1.364\times 10^{-8}$ $7.432\times 10^{-5}$ $6.299$ $1.696$ $-8.467$
TGARCH $9.072\times 10^{-9}$ $4.660\times 10^{-5}$ $1.102$ $0.711$ $-6.626$
HAR-RV $1.292\times 10^{-8}$ $5.798\times 10^{-5}$ $3.682$ $0.954$ $-8.412$
HAR-CJ $5.085\times 10^{-9}$ $3.364\times 10^{-5}$ $0.521$ $0.518$ $-8.700$
HAR-RS $4.453\times 10^{-9}$ $3.351\times 10^{-5}$ $0.840$ $0.580$ $-8.724$
HAR-REX $5.113\times 10^{-9}$ $3.444\times 10^{-5}$ $0.590$ $0.554$ $-8.696$
HAR-REQ $5.087\times 10^{-9}$ $3.353\times 10^{-5}$ $0.527$ $0.524$ $-8.697$
HAR-PD-RV $6.902\times 10^{-9}$ $3.966\times 10^{-5}$ $0.805$ $0.575$ $-8.606$
HAR-PD-CJ $5.062\times 10^{-9}$ $3.257\times 10^{-5}$ $\mathbf{0.445}$ $0.474$ $-8.699$
HAR-PD-RS $4.209\times 10^{-9}$ $3.460\times 10^{-5}$ $1.039$ $0.627$ $-8.659$
HAR-PD-REX $4.954\times 10^{-9}$ $3.294\times 10^{-5}$ $0.525$ $0.519$ $-8.700$
HAR-PD-REQ $4.888\times 10^{-9}$ $\mathbf{3.127\times 10^{-5}}$ $0.454$ $\mathbf{0.468}$ $-8.707$
LASSO-HAR-PD-CJ $5.072\times 10^{-9}$ $3.287\times 10^{-5}$ $0.464$ $0.484$ $-8.700$
LASSO-HAR-PD-REX $5.013\times 10^{-9}$ $3.325\times 10^{-5}$ $0.536$ $0.521$ $-8.698$
LASSO-HAR-PD-REQ $5.053\times 10^{-9}$ $3.473\times 10^{-5}$ $0.627$ $0.572$ $-8.692$
@@
Table 15: Out-of-sample forecast loss value of SSE for the future 22-step

### v2

@@ 4.5.1 Out-of-sample $R^{2}$ test
Here, $\mathrm{RV}_{t}$ represents the actual realized volatility, $\mathrm{RV}^{j}_{t}$ denotes the out-of-sample predicted value from the forecasting model, and $\mathrm{RV}^{0}_{t}$ refers to the out-of-sample predicted value from the benchmark model. If $R^{2}_{\mathrm{oos}}(\%)$ ¿ 0, it indicates that the model’s mean squared predictive error (MSPE-adjusted) is lower than that of the benchmark model, demonstrating positive predictive capability. If $R^{2}_{\mathrm{oos}}(\%)$ = 0, the forecasting performance of the model is identical to that of the benchmark model; if $R^{2}_{\mathrm{oos}}(\%)<0$ , the model performs worse than the benchmark model. The method for calculating the adjusted MSPE is detailed in Clark and West (2007).
@@ 4.5.2 Different rolling window lengths
GARCH 0.172-0.172 $3.881\times 10^{-9}$ $0.265$
EGARCH 0.633-0.633 $2.571\times 10^{-9}$ $0.852$
TGARCH 0.164-0.164 $4.219\times 10^{-9}$ $0.091$
GJRGARCH 0.136-0.136 $4.779\times 10^{-9}$ $0.226$
@@
LASSO-HAR-PD-RS $\mathbf{0.291}$ $3.331\times 10^{-8}$ $0.102$
LASSO-HAR-PD-REX $0.264$ $2.632\times 10^{-8}$ $0.065$
LASSO-HAR-PD-REQ $0.256$ $2.732\times 10^{-8}$ $0.064$
@@
As shown in Table 9, all models in the HAR-PD family exhibit an $R^{2}_{\mathrm{oos}}(\%)$ value significantly greater than zero, indicating their strong effectiveness in out-of-sample forecasting. Specifically, when comparing the HAR and HAR-PD model families, the LASSO-HAR-PD-RS model demonstrates the highest $R^{2}_{\mathrm{oos}}(\%)$ value in out-of-sample forecasting.
Furthermore, in Table 9, the $R^{2}_{\mathrm{oos}}(\%)$ values of the LASSO-HAR-PD model family and HAR-PD model family are also higher than those of the HAR model family, highlighting the overall improvement in predictive accuracy of the HAR-PD model family over traditional HAR model family. These findings further confirm that volatility forecasting models constructed from a path-dependence perspective exhibit superior predictive accuracy compared to benchmark models.
@@ 4.5.3 Different sample data
Table 10: Out-of-sample forecast loss value of CSI 300 for the one-step-ahead predictions
@@
GARCH $1.120\times 10^{-8}$ $5.216\times 10^{-5}$ $1.743$ $0.786$ 8.424-8.424
EGARCH $2.818\times 10^{-8}$ $7.593\times 10^{-5}$ $8.771$ $1.378$ 8.418-8.418
TGARCH $1.436\times 10^{-8}$ $6.703\times 10^{-5}$ $3.762$ $1.209$ 8.438-8.438
GJRGARCH $1.104\times 10^{-8}$ $5.122\times 10^{-5}$ $1.585$ $0.752$ 8.421-8.421
HAR-RV $1.228\times 10^{-8}$ $5.466\times 10^{-5}$ $3.409$ $0.865$ 8.364-8.364
HAR-CJ $8.565\times 10^{-9}$ $3.831\times 10^{-5}$ $0.477$ 0.496{0.496} 8.536-8.536
HAR-RS $7.919\times 10^{-9}$ $3.905\times 10^{-5}$ $1.003$ $0.571$ 8.553-8.553
HAR-REX $9.162\times 10^{-9}$ $4.421\times 10^{-5}$ $0.800$ $0.666$ 8.513-8.513
HAR-REQ $8.866\times 10^{-9}$ $\mathbf{3.818\times 10^{-5}}$ $\mathbf{0.433}$ $\mathbf{0.484}$ 8.525-8.525
HAR-PD-RV $9.513\times 10^{-9}$ $4.577\times 10^{-5}$ $1.241$ $0.731$ 8.501-8.501
hHAR-PD-CJ $8.422\times 10^{-9}$ $4.226\times 10^{-5}$ $0.688$ $0.604$ 8.529-8.529
HAR-PD-RS $7.774\times 10^{-9}$ $3.886\times 10^{-5}$ $1.771$ $0.574$ $\mathbf{-8.555}$
HAR-PD-REX $8.064\times 10^{-9}$ $4.062\times 10^{-5}$ $0.640$ $0.593$ 8.535-8.535
HAR-PD-REQ $\mathbf{7.740\times 10^{-9}}$ $3.956\times 10^{-5}$ $0.602$ $0.569$ 8.554-8.554
LASSO-HAR-PD-CJ $8.427\times 10^{-9}$ $4.250\times 10^{-5}$ $0.703$ $0.612$ 8.528-8.528
LASSO-HAR-PD-RS $7.638\times 10^{-9}$ $3.757\times 10^{-5}$ $0.956$ $0.590$ 8.597-8.597
LASSO-HAR-PD-REX $7.741\times 10^{-9}$ $3.951\times 10^{-5}$ $0.599$ $0.567$ 8.553-8.553
LASSSO-HAR-PD-REQ $8.283\times 10^{-9}$ $4.140\times 10^{-5}$ $0.647$ $0.600$ 8.526-8.526
@@
Table 10 shows that for one-step-ahead predictions, the HAR-REQ and HAR-PD-REQ models have lower loss values across two different loss functions. The HAR-PD-RS model achieves the lowest loss value under the QLIKE loss function. As shown in Table 16 for the five-step-ahead forecast and Table 17 for the twenty-two-step-ahead prediction, the proposed HAR-REQ model and its improved HAR-PD-REQ variant outperform other models in volatility forecasting under the MAE, HMSE, and HMAE loss functions. Table 18 presents the MCS test based on the above loss functions. The test results indicate that the HAR-PD model family and the LASSO-HAR-PD model family have passed the test many times.
@@ 5 Conclusion
In out-of-sample forecasting, the HAR-PD model family, constructed based on path dependence, consistently outperformed the original HAR model family, significantly improving prediction accuracy. This result remained significant across different datasets, indicating the strong robustness of the new model. Notably, under the rolling window and out-of-sample evaluation methods, the new model exhibited sustained superiority. It is worth mentioning that among all models examined in this study, the HAR-PD-RS model achieved the highest prediction accuracy in forecasting daily volatility, while the HAR-PD-REQ model demonstrated superior predictive capability in medium- and long-term (five-step-ahead prediction and twenty-two-step-ahead prediction) volatility forecasting.
@@ Appendix B Out-of-sample forecast loss value and MCS test
Table 11: Out-of-sample forecast loss value of SSE for the five-step-ahead predictions
Model MSE MAE HMSE HMAE QLIKE
GARCH $6.535\times 10^{-9}$ $4.592\times 10^{-5}$ $1.788$ $0.909$ 8.626-8.626
EGARCH $5.732\times 10^{-9}$ $5.463\times 10^{-5}$ $3.126$ $1.289$ 8.587-8.587
TGARCH $1.029\times 10^{-8}$ $5.463\times 10^{-5}$ $2.692$ $0.996$ 8.543-8.543
GJRGARCH $1.054\times 10^{-8}$ $5.387\times 10^{-5}$ $2.211$ $0.929$ 8.347-8.347
HAR-RV $7.068\times 10^{-9}$ $4.904\times 10^{-5}$ $1.936$ $1.005$ 8.608-8.608
HAR-CJ $6.220\times 10^{-9}$ $4.148\times 10^{-5}$ $1.013$ $0.746$ 8.647-8.647
HAR-RS $4.453\times 10^{-9}$ $3.351\times 10^{-5}$ $0.840$ $0.580$ 8.724-8.724
HAR-REX $5.113\times 10^{-9}$ $3.444\times 10^{-5}$ $0.590$ $0.554$ 8.696-8.696
HAR-REQ $5.087\times 10^{-9}$ $3.353\times 10^{-5}$ $0.527$ $0.524$ 8.697-8.697
HAR-PD-RV $6.902\times 10^{-9}$ $3.966\times 10^{-5}$ $0.805$ $0.575$ 8.606-8.606
HAR-PD-CJ $5.062\times 10^{-9}$ $3.257\times 10^{-5}$ $\mathbf{0.445}$ $0.474$ 8.699-8.699
HAR-PD-RS $4.209\times 10^{-9}$ $3.460\times 10^{-5}$ $1.039$ $0.627$ 8.659-8.659
HAR-PD-REX $4.954\times 10^{-9}$ $3.294\times 10^{-5}$ $0.525$ $0.519$ 8.700-8.700
HAR-PD-REQ $4.888\times 10^{-9}$ $\mathbf{3.127\times 10^{-5}}$ $0.454$ $\mathbf{0.468}$ 8.707-8.707
LASSO-HAR-PD-CJ $5.072\times 10^{-9}$ $3.287\times 10^{-5}$ $0.464$ $0.484$ 8.700-8.700
LASSO-HAR-PD-REX $5.013\times 10^{-9}$ $3.325\times 10^{-5}$ $0.536$ $0.521$ 8.698-8.698
LASSO-HAR-PD-REQ $5.053\times 10^{-9}$ $3.473\times 10^{-5}$ $0.627$ $0.572$ 8.692-8.692
  • 1.
Table 12: Out-of-sample forecast loss value of SSE for the twenty-two-step-ahead predictions
Model MSE MAE HMSE HMAE QLIKE
GARCH $7.191\times 10^{-9}$ $4.367\times 10^{-5}$ $1.349$ $0.824$ 8.619-8.619
EGARCH $9.048\times 10^{-9}$ $6.407\times 10^{-5}$ $4.796$ $1.562$ 8.378-8.378
TGARCH $9.774\times 10^{-9}$ $5.295\times 10^{-5}$ $0.728$ $0.706$ 7.679-7.679
GJRGARCH $1.023\times 10^{-8}$ $4.689\times 10^{-5}$ $0.763$ $0.607$ 8.149-8.149
HAR-RV $7.057\times 10^{-9}$ $4.871\times 10^{-5}$ $1.965$ $1.013$ 8.619-8.619
HAR-CJ $5.082\times 10^{-9}$ $3.311\times 10^{-5}$ $0.525$ $0.519$ 8.710-8.710
HAR-RS $4.413\times 10^{-9}$ $3.282\times 10^{-5}$ $0.846$ $0.578$ 8.736-8.736
HAR-REX $5.119\times 10^{-9}$ $3.396\times 10^{-5}$ $0.594$ $0.554$ 8.707-8.707
HAR-REQ $5.099\times 10^{-9}$ $3.308\times 10^{-5}$ $0.529$ $0.524$ 8.708-8.708
HAR-PD-RV $6.986\times 10^{-9}$ $3.955\times 10^{-5}$ $0.817$ $0.578$ 8.614-8.614
HAR-PD-CJ $5.062\times 10^{-9}$ $3.204\times 10^{-5}$ $0.446$ $0.474$ 8.710-8.710
HAR-PD-RS $\mathbf{4.213\times 10^{-9}}$ $3.422\times 10^{-5}$ $1.059$ $0.630$ 8.669-8.669
HAR-PD-REX $4.964\times 10^{-9}$ $3.253\times 10^{-5}$ $0.525$ $0.510$ 8.711-8.711
LASSO-HAR-PD-CJ $5.068\times 10^{-9}$ $3.232\times 10^{-5}$ $0.464$ $0.483$ 8.711{-8.711}
LASSO-HAR-PD-RS $\underline{4.377\times 10^{-9}}$ $\underline{\mathbf{3.097\times 10^{-5}}}$ $0.677$ $0.491$ 8.744-8.744
LASSO-HAR-PD-REX $5.020\times 10^{-9}$ $3.277\times 10^{-5}$ $0.535$ $0.519$ 8.709-8.709
LASSO-HAR-PD-REQ $5.081\times 10^{-9}$ $3.438\times 10^{-5}$ $0.632$ $0.574$ 8.702-8.702
  • 1.
Table 13: MCS test for the one-step-ahead volatility forecasting model of the SSE
@@
GARCH $0.000$ $0.000$ $0.029$ $0.029$ $0.034$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
EGARCH $0.000$ $0.000$ $0.067$ $0.012$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
TGARCH $0.000$ $0.000$ $0.994$ $0.090$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
GJRGARCH $0.000$ $0.000$ $0.822$ $0.029$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-RV $0.000$ $0.000$ $0.209$ $0.120$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-CJ $0.000$ $0.000$ $0.039$ $0.023$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-RS $1.000$ $0.004$ $1.000$ $0.233$ $1.000$ $0.002$ $1.000$ $0.181$ $1.000$ $0.000$ $0.893$ $0.165$
HAR-REX $0.128$ $0.308$ $1.000$ $0.626$ $0.000$ $0.000$ $1.000$ $0.482$ $0.000$ $0.000$ $0.918$ $0.460$
HAR-REQ $0.486$ $0.440$ $1.000$ $1.000$ $1.000$ $0.221$ $1.000$ $1.000$ $0.000$ $0.000$ $1.000$ $1.000$
HAR-PD-RV $0.000$ $0.000$ $1.000$ $0.233$ $0.000$ $0.000$ $0.107$ $0.098$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-PD-CJ $1.000$ $1.000$ $1.000$ $1.000$ $0.000$ $1.000$ $0.735$ $1.000$ $1.000$ $0.992$ $0.499$ $1.000$
HAR-PD-RS $0.694$ $0.004$ $1.000$ $1.000$ $0.427$ $0.437$ $1.000$ $1.000$ $0.340$ $0.399$ $1.000$ $1.000$
HAR-PD-REX $0.994$ $1.000$ $1.000$ $\mathbf{1.000}$ $0.802$ $0.656$ $\mathbf{1.000}$ $\mathbf{1.000}$ $0.678$ $0.607$ $\mathbf{1.000}$ $\mathbf{1.000}$
HAR-PD-REQ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$
LASSO-HAR-PD-CJ $0.672$ $0.660$ $1.000$ $1.000$ $0.380$ $0.547$ $1.000$ $0.975$ $0.329$ $0.399$ $0.998$ $0.977$
@@
Table 14: MCS test for the five-step-ahead volatility forecasting model of the SSE
@@
TGARCH $0.000$ $0.000$ $0.046$ $0.072$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
GJRGARCH $0.000$ $0.000$ $0.964$ $0.029$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.043$ $0.083$
HAR-RV $0.000$ $0.000$ $0.298$ $0.080$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
HAR-CJ $0.251$ $0.994$ $1.000$ $0.469$ $0.000$ $0.000$ $1.000$ $0.423$ $0.192$ $0.659$ $0.993$ $0.486$
HAR-RS $1.000$ $0.008$ $1.000$ $0.098$ $1.000$ $0.000$ $1.000$ $0.085$ $1.000$ $0.002$ $0.998$ $0.096$
HAR-REX $0.129$ $0.647$ $1.000$ $0.530$ $0.000$ $0.000$ $1.000$ $0.423$ $0.084$ $0.659$ $1.000$ $0.602$
HAR-REQ $0.664$ $1.000$ $1.000$ $1.000$ $0.237$ $0.640$ $1.000$ $1.000$ $0.403$ $0.779$ $1.000$ $0.931$
HAR-PD-RV $0.000$ $0.000$ $0.246$ $0.469$ $0.000$ $0.000$ $0.107$ $0.097$ $0.000$ $0.000$ $0.117$ $0.118$
HAR-PD-CJ $0.882$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $0.802$ $1.000$ $0.717$ $0.805$ $1.000$ $0.724$
HAR-PD-RS $0.943$ $0.005$ $1.000$ $1.000$ $0.665$ $0.640$ $1.000$ $1.000$ $0.989$ $0.001$ $1.000$ $0.997$
HAR-PD-REX $1.000$ $1.000$ $1.000$ 1.000 $1.000$ 1.000 $1.000$ 1.000 $0.995$ $0.998$ 1.000 1.000
HAR-PD-REQ $1.000$ $1.000$ 1.000 $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $1.000$ $0.989$
LASSO-HAR-PD-CJ $0.794$ $1.000$ $1.000$ $1.000$ $0.345$ $0.995$ $0.971$ $0.440$ $0.654$ $0.781$ $0.995$ $0.646$
@@
Table 15: MCS test for the twenty-two-step-ahead volatility forecasting model of the SSE
@@
Table 16: Out-of-sample forecast loss value of CSI 300 for the five-step-ahead predictions
@@
GARCH $1.176\times 10^{-8}$ $5.154\times 10^{-5}$ $1.584$ $0.746$ 8.392-8.392
EGARCH $1.651\times 10^{-8}$ $6.938\times 10^{-5}$ $3.995$ $1.209$ 8.339-8.339
TGARCH $1.027\times 10^{-8}$ $4.512\times 10^{-5}$ $0.892$ $0.647$ 8.483-8.483
GJRGARCH $1.133\times 10^{-8}$ $4.905\times 10^{-5}$ $1.247$ $0.673$ 8.403-8.403
HAR-RV $1.286\times 10^{-8}$ $5.516\times 10^{-5}$ $3.669$ $0.878$ 8.367-8.367
HAR-CJ $9.018\times 10^{-9}$ $3.846\times 10^{-5}$ $0.476$ $0.497$ 8.543-8.543
HAR-RS $7.581\times 10^{-9}$ $\mathbf{3.795\times 10^{-5}}$ $0.832$ $0.552$ 8.514-8.514
HAR-REX $9.645\times 10^{-9}$ $4.461\times 10^{-5}$ $0.807$ $0.671$ 8.519-8.519
HAR-REQ $9.352\times 10^{-9}$ $\underline{3.840\times 10^{-5}}$ $\underline{\mathbf{0.433}}$ $\underline{\mathbf{0.485}}$ 8.531-8.531
HAR-PD-RV $9.985\times 10^{-9}$ $4.602\times 10^{-5}$ $1.257$ $0.736$ 8.507-8.507
HAR-PD-CJ $8.873\times 10^{-9}$ $4.258\times 10^{-5}$ $0.696$ $0.611$ 8.533-8.533
HAR-PD-RS $\underline{\mathbf{7.575\times 10^{-9}}}$ $3.907\times 10^{-5}$ $0.948$ $0.583$ 8.541-8.541
HAR-PD-REX $8.502\times 10^{-9}$ $4.152\times 10^{-5}$ $0.662$ $0.607$ 8.540-8.540
LASSO-HAR-PD-CJ $8.878\times 10^{-9}$ $4.282\times 10^{-5}$ $0.712$ $0.619$ 8.533-8.533
LASSO-HAR-PD-RS $7.576\times 10^{-9}$ $3.907\times 10^{-5}$ $0.948$ $0.581$ 8.542-8.542
LASSO-HAR-PD-REX $8.745\times 10^{-9}$ $4.247\times 10^{-5}$ $0.682$ $0.618$ 8.531-8.531
LASSO-HAR-PD-REQ $8.182\times 10^{-9}$ $4.034\times 10^{-5}$ $0.629$ $0.583$ 8.558-8.558
@@
Table 17: Out-of-sample forecast loss value of CSI 300 for the twenty-two-step-ahead predictions